Consider the system of linear equations below.
3x+y=-4
-5x+y=8
A. Graph the system of equations on the coordinate plane. Label each line on your graph and show any work.
B. Use your graph from Part A to estimate the solution for the system of equations. Explain how you determined your estimate.

Answer by mananth(16946)

3 x + 1 y = -4 .............1
-5 x + 1 y = 8 .............2
Plug x=0,& y=0 in both the equations to get the y & x intercepts of the equations
equation1
when x= 0, y= -4 ( 0 , -4 )
when y = 0, x= -4/3 ( -4/3 , 0 )
Equation 2
when x= 0, y= 8 ( 0 , 8 )
when y = 0, x= -8/5 ( -8/5 , 0 )
Plot the points ,draw the lines.
The point of intersection is the solution

The point of intersection is (-1.5,0.5)